# Local Level Dynamic Random Partition Models for Changepoint Detection

### v1

@@ Abstract
Motivated by an increasing demand for models that can effectively describe features of complex multivariate time series, e.g. from sensor data in biomechanics, motion analysis, and sports science, we introduce a novel state-space modeling framework where the state equation encodes the evolution of latent partitions of the data over time. Building on the principles of dynamic linear models, our approach develops a random partition model capable of linking data partitions to previous ones over time, using a straightforward Markov structure that accounts for temporal persistence and facilitates changepoint detection. The selection of changepoints involves multiple dependent decisions, and we address this time-dependence by adopting a non-marginal false discovery rate control. This leads to a simple decision rule that ensures more stringent control of the false discovery rate compared to approaches that do not consider dependence. The method is efficiently implemented using a Gibbs sampling algorithm, leading to a straightforward approach compared to existing methods for dependent random partition models. Additionally, we show how the proposed method can be adapted to handle multi-view clustering scenarios. Simulation studies and the analysis of a human gesture phase dataset collected through various sensing technologies show the effectiveness of the method in clustering multivariate time series and detecting changepoints.
@@ 1 Introduction
Let $\bm{Y}_{1},\ldots,\bm{Y}_{T}$ denote a multivariate time series, where each $\bm{Y}_{t}$ is a $n$-dimensional vector, i.e., $\bm{Y}_{t}=\{Y_{1,t},Y_{2,t},\ldots,Y_{n,t}\}$, observed on $n$ units over $T$ time points. To illustrate, consider our application in Section 5, where we examine scalar velocity data obtained from acceleration sensor units placed on the hands and wrists of a subject. Dynamic linear models (DLMs) are commonly employed for the analysis of time-series data due to their flexibility and adaptability in handling diverse situations (Petris et al., 2009). They define a class of state-space models and are characterized by a system of two equations: an observation equation, which describes the observed data as a linear combination of latent state variables with noise, and a state equation that describes how latent states evolve over time, thereby tracking the underlying dynamics of the system. We introduce our contribution by referring to a simple yet fundamental DLM, the local level model, which describes the observed data as composed of a level component plus random noise,
$Y_{i,t}=\beta_{i,t}+\varepsilon_{i},$ (1)
@@
where $\varepsilon_{i}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{N}(0,\tau^{2})$, with $\text{N}(\mu,\sigma^{2})$ denoting the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. The vector $\bm{\beta}_{t}=(\beta_{1,t},\ldots,\beta_{n,t})$ represents the underlying level or trend of the time series at time $t$. In a typical local level model, the evolution of $\bm{\beta}_{t}$ over time is modeled as a random walk, i.e., the level at time $t$ is predicted to be the same as the level at time $t-1$, plus some random noise. If the variance of the random noise is small, this assumption implies smooth processes over time. Despite its simplicity, the local level model illustrates the fundamental characteristics of many time-series models, and will serve as a basic example throughout.
@@
One of the main goals of our biometric data application is to study how measurements from different units cluster or group over time, as well as how these clusters evolve during an activity, to understand how different body parts cooperate during different stages of a gesture or movement. Furthermore, the detection of changepoints in the clustering of body parts between gesture stages can reveal important insights into the motion and inform gesture detection algorithms. In the Bayesian setting, Bayesian nonparametric (BNP) models are a popular choice for clustering dynamic behavior over time (Quintana et al., 2022). BNP models do not require the upfront specification of the number of clusters; instead, they allow for posterior inference on cluster allocations directly from the data. Existing BNP approaches for clustering time series data vary in terms of motivation, application, and how time-dependence is introduced. For example, some approaches build on the stick-breaking representation of the Dirichlet Process (Ferguson, 1973; Sethuraman, 1994). In this context, Antoniano-Villalobos and Walker (2016) have developed a stationary Markov model where both the transition and stationary densities are nonparametric infinite mixture models. Nieto-Barajas and Contreras-Cristán (2014) have clustered temporal data while considering several features typical of time series data (e.g., trends and seasonality). BNP autoregressive model are discussed in Kalli and Griffin (2018), De Iorio et al. (2019), DeYoreo and Kottas (2018), and Beraha et al. (2022), among others. Alternatively, other authors have explored generalizations of the Pólya urn scheme of Blackwell et al. (1973), see, e.g. Caron et al. (2007), Caron et al. (2017) and Cassese et al. (2019).
@@
A solution to address this challenge involves directly modeling the sequence of random partitions. In this context, Page et al. (2022) introduced a dependent random partition model (Hartigan, 1990; Barry and Hartigan, 1992) that incorporates auxiliary variables that determine whether a unit, for which the cluster allocation at time $t-1$ has been determined, should be considered for possible cluster reallocation at time $t$. As a result, a change of partitions over time is obtained as a by-product of individual clustering allocations. This may lead to a high number of cluster configurations and may hinder the identification of partition changepoints through time, which is crucial, for instance, in identifying gesture phases. Quinlan et al. (2022) presented a method that aims to correlate partitions with the detection of changepoints in multivariate time series in the context of contagion phenomena. However, following Martinez and Mena (2014), they specify only a single unique partition of contiguous clusters for each time series. The focus of their model is on multivariate changepoint detection rather than data partitioning, with changepoints occurring at different time points for each time series.
@@
Recent methods have been proposed for multiview clustering that allow for dependence between several experimental conditions or features. For example, Franzolini et al. (2023) employ a conditional partially exchangeable model to induce dependence between clustering configurations of the same subjects across different features. Similarly, Dombowsky and Dunson (2023) develop a nonparametric prior that generates dependent random distributions by centering a Dirichlet process on a random product measure, with a single parameter controlling the dependence. Our approach can also be adapted for multi-view clustering scenarios. By incorporating a latent-state representation of the partitions, our framework builds a hierarchical model that effectively captures the dependence between partitions, describing changes across different groups or experimental conditions. Additionally, we show that the dependence between the partitions generated by the hierarchical model for multiview clustering in Dombowsky and Dunson (2023) coincides with that one of a reparametrization of our prior specification. Finally, we should note that while constructing the prior distribution for the partition bears some resemblance to the use of spike-and-slab priors for variable selection (Tadesse and Vannucci, 2021), dealing with partitions introduces additional complexity in the motivation, modeling and computation.
The article is organized as follows: in Section 2, we present our model and discuss its main properties. Section 2.2 discusses the alternative multiview formulation of our model. In Section 3 we cover posterior inference, including computational methods and decision theory-based changepoint detection. In Section 4 we describe two simulation studies that highlight the key aspects of our model. In Section 5, we present an application to the analysis of human gesture data, and finally, in Section 6, we provide concluding remarks and outline future directions. Proofs and additional details on computations and illustrations are available in the Supplementary Material.
@@ 2.1 Prior properties of the PSM
with $-\infty\leq\sigma<1$, and where, for any $n\geq 1$ and any $k=1,\ldots,n$, the set of non-negative weights $V_{n,k}$ satisfies the following recursive equation $V_{n,k}=(n-\sigma k)V_{n+1,k}+V_{n+1,k+1}$, and is such that $V_{1,1}=1$. Notable elements of the class of Gibbs-type priors are the Dirichlet and the Pitman-Yor processes (Ferguson, 1973; Pitman and Yor, 1997), mixtures of symmetric Dirichlet distributions (Gnedin and Pitman, 2005), the normalized inverse Gaussian processes (Lijoi et al., 2005) and the normalized generalized gamma process (Lijoi et al., 2007). Throughout the paper, as a running example for the base partition distribution $p^{*}(\cdot)$, we will consider the Chinese restaurant process (CRP, Pitman, 2002) with concentration parameter $\theta>0$, denoted as $p_{\text{CRP}}(\cdot)$, which can be obtained as the distribution on the partitions implied by the Dirichlet process. The corresponding EPPF is readily obtained by setting $\sigma=0$ and $V_{n,|\pi_{t}|}=\theta^{|\pi_{t}|}/(\theta)_{n}$ in (6), where $(\theta)_{n}=\theta(\theta+1)\cdots(\theta+n-1)$ denotes the ascending factorial. For each specification within the class of Gibbs-type priors, the weights $V_{n,k}$ can be computed recursively or estimated via Monte Carlo (Arbel et al., 2017).
@@
$\varphi_{t_{1},t_{2}}=\mathbbm{E}[R(\pi_{t_{1}},\pi_{t_{2}})]=\sum_{\pi_{t_{1}},\pi_{t_{2}}\in\mathscr{P}}R(\pi_{t_{1}},\pi_{t_{2}})\,p(\pi_{t_{1}},\pi_{t_{2}}),$
@@
When $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$, the ERI for two random partitions at subsequent times modeled according to the proposed PSM is obtained by substituting the appropriate specification of $V_{2,2}$ in Proposition 1, which leads to the following corollary.
@@
Suppose $t_{2}=t_{1}+1$ and $\eta_{t_{1}}=\eta_{t_{2}}=\eta$ and let $p^{\ast}(\cdot)=p_{\text{CRP}}(\cdot)$. Then the ERI for two random partitions $\pi_{t_{1}}$ and $\pi_{t_{2}}$ modeled according to a PSM is given by
$\varphi_{t_{1},t_{2}}=1-\frac{2\theta}{(\theta+1)^{2}}\eta.$ (8)
The expression for the ERI in (8) coincides, up to a reparametrization of the dependence parameter $\eta$, with an analogous result derived in Dombowsky and Dunson (2023). Interestingly, the ERI approaches one both when $\theta\to 0$ and when $\theta\to\infty$. This behavior is not surprising as, when $\theta$ approaches zero, the distribution of both random partitions tend to assign mass one to the partition with only one block. Similarly, when $\theta$ increases, the two distributions accumulate mass at the partition composed of $n$ blocks of size one. In both cases, the two random partitions tend to take the same configuration, regardless of the value taken by the dependence parameter $\eta$. Also in this case, (8) can be written as a convex linear combination of the ERI of two identical random partitions, that is one, and the ERI of two independent random partitions both distributed as a Chinese restaurant process, that is $(\theta^{2}+1)/(\theta+1)^{2}$. The next proposition describes how the dependence between two lagged random partitions modeled through the proposed PSM changes as a function of the lag $t_{2}-t_{1}$.
@@
The ERI is thus a decreasing function of the time lag $t_{2}-t_{1}$. Furthermore, regardless of the value of the dependence parameter $\eta$, the limit of $\varphi_{t_{1},t_{2}}$ as $(t_{2}-t_{1})\to\infty$ tends towards the ERI of two independent random partitions with identical distribution implied by a Gibbs-type prior. The next corollary specializes the previous result to the case $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$.
@@
Suppose $t_{1}<t_{2}$ and let $\eta_{t_{1}}=\eta_{t_{2}}=\eta$. If $p^{\ast}(\cdot)=p_{\text{CRP}}(\cdot)$, then the ERI for two random partitions $\pi_{t_{1}}$ and $\pi_{t_{2}}$ modeled according to a PSM is given by
$\varphi_{t_{1},t_{2}}=1-\frac{2\theta}{(\theta+1)^{2}}[1-(1-\eta)^{t_{2}-t_{1}}].$ (9)
@@ 2.2 Multiview clustering representation of the LLDPM
In many real-world applications, data can be collected from various sources or represented using different feature sets, creating multiple views of the same underlying entities. In this section, we present an alternative representation of our model that establishes a hierarchical structure in the dependence of partitions, deviating from the assumption of temporally dependent partitions where each partition depends on the previous one. This hierarchical representation effectively describes changes in the partition across different groups or experimental conditions, offering a flexible and interpretable approach to modeling and understanding partition relationships. Such representation exemplifies a multiview clustering model, bearing close connections to the partition introduced by Dombowsky and Dunson (2023) and related to the multiview representation in Franzolini et al. (2023). As noted below (8), the same expected dependence between partitions induced by the hierarchical model for multiview clustering in Dombowsky and Dunson (2023) can be achieved by our model by reparametrizing the dependence parameter $\eta$. In their prior specification, dependence is introduced through the concentration parameter of a Dirichlet process, which determines the tendency of data points to form distinct groups. The main difference between our model and theirs lies in the way the dependence is induced. While their models specify a dependence structure on the atoms within a hierarchical model on a product space, our model introduces dependence directly at the level of partitions, providing a more streamlined approach to modeling partition relationships.
We begin by examining the joint distribution of two subsequent random partitions in the PSM we proposed, where $T$ can now be interpreted as the number of different views, groups or experimental conditions, depending on the context of the application. Without loss of generality, we consider the case where $T=2$ and thus focus on the joint distribution of $\pi_{1}$ and $\pi_{2}$, as implied by the PSM, which is
@@
$\begin{split}\tilde{\pi}&\,\sim\,p^{*}(\tilde{\pi}).\end{split}$ (11)
@@
where $\tilde{\gamma}_{t}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{Bern}(\tilde{\eta})$, $t=1,2$, controlling the degree to which $\pi_{1}$ and $\pi_{2}$ can differ from the parent partition $\tilde{\pi}$. A higher value of $\tilde{\eta}$ increases the chances of $\pi_{t}$ being different from $\tilde{\pi}$, allowing for more flexibility and individual variations in the partition structure. Conversely, a lower value of $\tilde{\eta}$ makes it more likely that $\pi_{t}$ will follow the same partition structure as $\tilde{\pi}$, enforcing a stronger dependence between the two. Note that, in model (11)–(12), the distribution of $\pi_{2}$ is independent of that of $\pi_{1}$, given the partition $\tilde{\pi}$. The joint distribution of the random partitions $\pi_{1}$ and $\pi_{2}$, the common parent partition $\tilde{\pi}$, and the dependence variables $(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$, here denoted by $p(\bm{\pi}_{1:2},\tilde{\pi},\tilde{\gamma}_{1},\tilde{\gamma}_{2})$, can thus be written as
@@
Hence, the distribution of $\pi_{1}$ and $\pi_{2}$ is obtained by marginalizing the last expression with respect to $\tilde{\pi}$ and $(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$, which gives (10), provided that $\tilde{\eta}=1-\sqrt{1-\eta}$.
Hence, the local level partition model’s joint distribution on the partitions can be viewed as a special case of a hierarchical partition model, when considered marginally. This representation highlights the connection between the time-varying formulation of the PSM, as specified through (3)-(4), and an exchangeable model on the partitions, under the assumptions of identical marginal distributions of the partitions, $p^{*}(\pi_{t})$, and equal Bernoulli probabilities $\tilde{\eta}$ across groups. This hierarchical representation can be extended to accommodate $T\geq 2$ groups, with one parent random partition driving the distribution of $T$ identically distributed random partitions. Moreover, the assumption of a common Bernoulli probability $\tilde{\eta}$ across groups can be relaxed to accommodate for group-specific parameters $\tilde{\eta}_{t}$, with $t=1,\ldots,T$. Expressing the local level partition model as a hierarchical model provides an alternative viewpoint to appreciate its flexibility in capturing both similarities and dissimilarities among the partitions. The hierarchical structure allows for the sharing of information across groups while still accommodating individual variations in the partition structures.

### v2

@@ Abstract
Motivated by an increasing demand for models that can effectively describe features of complex multivariate time series, e.g. from sensor data in biomechanics, motion analysis, and sports science, we introduce a novel state-space modeling framework where the state equation encodes the evolution of latent partitions of the data over time. Building on the principles of dynamic linear models, our approach develops a random partition model capable of linking data partitions to previous ones over time, using a straightforward Markov structure that accounts for temporal persistence and facilitates changepoint detection. The selection of changepoints involves multiple dependent decisions, and we address this time-dependence by adopting a non-marginal false discovery rate control. This leads to a simple decision rule that ensures more stringent control of the false discovery rate compared to approaches that do not consider dependence. The method is efficiently implemented using a Gibbs sampling algorithm, leading to a straightforward approach compared to existing methods for dependent random partition models. Additionally, we show how the proposed method can be adapted to handle multiview clustering scenarios. Simulation studies and the analysis of a human gesture phase dataset collected through various sensing technologies show the effectiveness of the method in dynamically clustering multivariate time series and detecting changepoints.
@@ 1 Introduction
Let $\bm{Y}_{1:T}=\{\bm{Y}_{1},\ldots,\bm{Y}_{T}\}$ denote a multivariate time series, where each $\bm{Y}_{t}$ is a $n$-dimensional vector, i.e., $\bm{Y}_{t}=(Y_{1,t},Y_{2,t},\ldots,Y_{n,t})$, observed on $n$ units over $T$ time points. To illustrate, consider our application in Section 5.1, where we examine scalar velocity data obtained from acceleration sensor units placed on the hands and wrists of a subject. dynamic linear models (DLMs) are commonly employed for the analysis of time-series data due to their flexibility and adaptability in handling diverse situations (Petris et al., 2009). They define a class of state-space models and are characterized by a system of two equations: an observation equation, which describes the observed data as a linear combination of latent state variables with noise, and a state equation that describes how latent states evolve over time, thereby tracking the underlying dynamics of the system. We introduce our contribution by referring to a simple yet fundamental DLM, the local level model, which describes the observed data as composed of a level component plus random noise,
$Y_{i,t}=\beta_{i,t}+\varepsilon_{i,t},$ (1)
@@
where $\varepsilon_{i,t}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{N}(0,\tau^{2})$, with $\text{N}(\mu,\sigma^{2})$ denoting the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. The vector $\bm{\beta}_{t}=(\beta_{1,t},\ldots,\beta_{n,t})$ represents the underlying level or trend of the time series at time $t$. In a typical local level model, the evolution of $\bm{\beta}_{t}$ over time is modeled as a random walk, i.e., the level at time $t$ is predicted to be the same as the level at time $t-1$, plus some random noise. If the variance of the random noise is small, this assumption implies smooth processes over time. Despite its simplicity, the local level model illustrates the fundamental characteristics of many time-series models, and will serve as a basic example throughout.
@@
One of the main goals of our biometric data application is to study how measurements from different units cluster or group over time, as well as how these clusters evolve during an activity, to understand how different body parts cooperate during different stages of a gesture or movement. Furthermore, the detection of changepoints in the clustering of body parts between gesture stages can reveal important insights into the motion and inform gesture detection algorithms. In the Bayesian setting, Bayesian nonparametric (BNP) models are a popular choice for clustering dynamic behavior over time (Quintana et al., 2022). BNP models do not require the upfront specification of the number of clusters; instead, they allow for posterior inference on cluster allocations directly from the data. Existing BNP approaches for clustering time series data vary in terms of motivation, application, and how time-dependence is introduced. For example, some approaches build on the stick-breaking representation of the Dirichlet Process (Ferguson, 1973; Sethuraman, 1994). In this context, Antoniano-Villalobos and Walker (2016) have developed a stationary Markov model where both the transition and stationary densities are nonparametric infinite mixture models. Nieto-Barajas and Contreras-Cristán (2014) have clustered temporal data while considering several features typical of time series data (e.g., trends and seasonality). BNP autoregressive model are discussed in Kalli and Griffin (2018), De Iorio et al. (2023), DeYoreo and Kottas (2018), Pavani and Quintana (2024), and Beraha et al. (2023), among others. Alternatively, other authors have explored generalizations of the Pólya urn scheme of Blackwell et al. (1973), see, e.g. Caron et al. (2007), Caron et al. (2017) and Cassese et al. (2019).
@@
An alternative approach involves directly modeling the sequence of random partitions. In this context, Page et al. (2022) introduced a dependent random partition model (Hartigan, 1990; Barry and Hartigan, 1992) that incorporates auxiliary variables that determine whether a unit, for which the cluster allocation at time $t-1$ has been determined, should be considered for possible cluster reallocation at time $t$.As a result, changes in partitions over time are obtained as a by-product of individual clustering assignments. However, this can lead to a high number of changes in cluster configurations and may hinder the detection of partition changepoints over time, which is crucial, for instance, in identifying gesture phases. Paganin et al. (2023) proposed a model for a sequence of random partitions over time, considering individual allocation probabilities for each unit rather than the whole partition at each time point. Cremaschi et al. (2023) introduced a Bayesian changepoint model for spatiotemporal data, where partition changes in areal units over time are governed by a spatial association parameter. Their model employs a random partition prior that incorporates spatial features, promoting the co-clustering of nearby areal units based on proximity. Quinlan et al. (2022) presented a method that aims to correlate partitions with the detection of changepoints in multivariate time series in the context of contagion phenomena. However, following Martinez and Mena (2014), they specify only a single unique partition of contiguous clusters for each time series. The focus of their model is on multivariate changepoint detection rather than data partitioning, with changepoints occurring at different time points for each time series. Recently, Corradin et al. (2024) proposed a Bayesian model for clustering time series, where two observations are assigned to the same group if they exhibit structural changes in their behavior at the same time. Their approach assumes that units within the same cluster share changepoints across all time points, with no possibility of shifts of units in different clusters at different time points.
@@
Recent methods have been proposed for multiview clustering that allow for dependence between several experimental conditions or features. For example, Franzolini et al. (2023) employ a conditional partially exchangeable model to induce dependence between clustering configurations of the same subjects across different features. Similarly, Dombowsky and Dunson (2023) develop a nonparametric prior that generates dependent random distributions by centering a Dirichlet process on a random product measure, with a single parameter controlling the dependence. Our approach can also be adapted for multiview clustering scenarios. By incorporating a latent-state representation of the partitions, our framework builds a hierarchical model that effectively captures the dependence between partitions, describing changes across different groups or experimental conditions. Additionally, we show that the dependence between the partitions generated by the hierarchical model for multiview clustering in Dombowsky and Dunson (2023) coincides with that one of a reparametrization of our prior specification. Finally, we should note that while constructing the prior distribution for the partition bears some resemblance to the use of spike-and-slab priors for variable selection (Tadesse and Vannucci, 2021), dealing with partitions introduces additional complexity in the motivation, modeling and computation.
The article is organized as follows: in Section 2, we present our model and discuss its main properties. Section 2.2 discusses the alternative multiview formulation of our model. In Section 3 we cover posterior inference, including computational methods and decision theory-based changepoint detection. In Section 4 we describe two simulation studies that highlight the key aspects of our model. In Section 5, we present an application to the analysis of human gesture data, and illustrate its performance in two-view clustering using data from a collaborative perinatal project. Finally, in Section 6, we provide concluding remarks and outline future directions. Proofs and additional details on computations and illustrations are available in the Supplementary Material.
@@ 2.1 Prior properties of the PSM
with $-\infty\leq\sigma<1$, and where, for any $n\geq 1$ and any $k=1,\ldots,n$, the set of non-negative weights $V_{n,k}$ satisfies the following recursive equation $V_{n,k}=(n-\sigma k)V_{n+1,k}+V_{n+1,k+1}$, and is such that $V_{1,1}=1$. Notable elements of the class of Gibbs-type priors are the Dirichlet and the Pitman–Yor processes (Ferguson, 1973; Pitman and Yor, 1997), mixtures of symmetric Dirichlet distributions (Gnedin and Pitman, 2005), the normalized inverse Gaussian processes (Lijoi et al., 2005) and the normalized generalized gamma process (Lijoi et al., 2007). Throughout the paper, we use the Chinese restaurant process (CRP, Pitman, 2006) and its generalization, the two-parameter CRP (2-CRP, Pitman, 2006), as running examples for the base partition distribution $p^{*}(\cdot)$. The CRP, denoted $p_{\text{CRP}}(\cdot)$, is parametrized by a concentration parameter $\theta>0$ and corresponds to distribution on the partitions induced by the Dirichlet process. The corresponding EPPF is readily obtained by setting $\sigma=0$ and $V_{n,|\pi_{t}|}=\theta^{|\pi_{t}|}/(\theta)_{n}$ in (6), where $(\theta)_{n}=\theta(\theta+1)\cdots(\theta+n-1)$ denotes the ascending factorial. The 2-CRP, denoted $p_{\sigma}(\cdot)$, introduces an additional discount parameter $\sigma\in[0,1)$, with $\theta>-\sigma$, and coincides with the distribution on the partitions associated with the Pitman–Yor process (Perman et al., 1992; Pitman and Yor, 1997). The corresponding EPPF is obtained by setting $V_{n,|\pi_{t}|}=\prod_{j=1}^{|\pi_{t}|-1}(\theta+j\sigma)/(\theta+1)_{n-1}$ in (6). For each specification within the class of Gibbs-type priors, the weights $V_{n,k}$ can be computed recursively or estimated via Monte Carlo (Arbel et al., 2017).
@@
$\varphi_{t_{1},t_{2}}=\mathbb{E}[R(\pi_{t_{1}},\pi_{t_{2}})]=\sum_{\pi_{t_{1}},\pi_{t_{2}}\in\mathscr{P}}R(\pi_{t_{1}},\pi_{t_{2}})\,p(\pi_{t_{1}},\pi_{t_{2}}),$
@@
When $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$ or $p^{*}(\cdot)=p_{\sigma}(\cdot)$, the ERI for two random partitions at subsequent times modeled according to the proposed PSM is obtained by substituting the appropriate specification of $V_{2,2}$ in Proposition 1, which leads to the following corollary.
@@
Suppose $t_{2}=t_{1}+1$ and $\eta_{t_{1}}=\eta_{t_{2}}=\eta$. Then the ERI for two random partitions $\pi_{t_{1}}$ and $\pi_{t_{2}}$ modeled according to a PSM is given by
$\varphi_{t_{1},t_{2}}$ $=1-\frac{2\theta}{(\theta+1)^{2}}\eta$ if $\text{if }p^{*}(\cdot)=p_{\textsc{CRP}}(\cdot),$ (8)
$\varphi_{t_{1},t_{2}}$ $=1-\ \frac{2(\theta+\sigma)(1-\sigma)}{(\theta+1)^{2}}\eta$ if $\text{if }p^{*}(\cdot)=p_{\sigma}(\cdot).$ (9)
The expression for the ERI in (8) coincides, up to a reparametrization of the dependence parameter $\eta$, with an analogous result derived in Dombowsky and Dunson (2023). As expected, the ERI in (8) is recovered by setting $\sigma=0$ in (9). Interestingly, the ERI approaches one both when $\theta\to-\sigma$ and when $\theta\to\infty$. This behavior is not surprising as, when $\theta$ approaches $-\sigma$, the distribution of both random partitions tend to assign mass one to the partition with only one block. Similarly, when $\theta$ increases, the two distributions accumulate mass at the partition composed of $n$ blocks of size one. In both cases, the two random partitions tend to take the same configuration, regardless of the value taken by the dependence parameter $\eta$. Also in this case, (8) and (9) can be written as convex linear combinations of the ERI of two identical random partitions, that is one, and the ERI of two independent and identically distributed random partitions, that is $(\theta^{2}+1)/(\theta+1)^{2}$ if $p^{*}(\cdot)=p_{\textsc{CRP}}(\cdot)$ and $(\theta^{2}+1)/(\theta+1)^{2}+2\sigma(\theta+\sigma-1)/(\theta+1)^{2}$ if $p^{*}(\cdot)=p_{\sigma}(\cdot)$.
The next proposition describes how the dependence between two lagged random partitions modeled through the proposed PSM changes as a function of the lag $t_{2}-t_{1}$.
@@
The ERI is thus a decreasing function of the time lag $t_{2}-t_{1}$. Furthermore, regardless of the value of the dependence parameter $\eta$, the limit of $\varphi_{t_{1},t_{2}}$ as $(t_{2}-t_{1})\to\infty$ tends towards the ERI of two independent random partitions with identical distribution implied by a Gibbs-type prior. The next corollary specializes the previous result to the cases $p^{*}(\cdot)$ being a CRP or a 2-CRP.
@@
Suppose $t_{1}<t_{2}$ and let $\eta_{t_{1}}=\eta_{t_{2}}=\eta$. Then the ERI for two random partitions $\pi_{t_{1}}$ and $\pi_{t_{2}}$ modeled according to a PSM is given by
$\varphi_{t_{1},t_{2}}$ $=1-\frac{2\theta}{(\theta+1)^{2}}[1-(1-\eta)^{t_{2}-t_{1}}]$ if $\text{if }p^{*}(\cdot)=p_{\textsc{CRP}}(\cdot),$ (10)
$\varphi_{t_{1},t_{2}}$ $=1-\frac{2(\theta+\sigma)(1-\sigma)}{(\theta+1)^{2}}[1-(1-\eta)^{t_{2}-t_{1}}]$ if $\text{if }p^{*}(\cdot)=p_{\sigma}(\cdot).$ (11)
@@ 2.2 Multiview clustering representation of the LLDPM
In many real-world applications, data can be collected from various sources or represented using different feature sets, creating multiple views of the same underlying entities, with each view defined on its own support. In this section, we present an alternative representation of our model that establishes a hierarchical structure in the dependence of partitions, deviating from the assumption of temporally dependent partitions where each partition depends on the previous one. This hierarchical representation effectively describes changes in the partition across different groups or experimental conditions, offering a flexible and interpretable approach to modeling and understanding partition relationships. Such representation exemplifies a multiview clustering model, bearing close connections to the partition introduced by Dombowsky and Dunson (2023) and related to the multiview representation in Franzolini et al. (2023). As noted below (8), the same expected dependence between partitions induced by the hierarchical model for multiview clustering in Dombowsky and Dunson (2023) can be achieved by our model by reparametrizing the dependence parameter $\eta$. In their prior specification, dependence is introduced through the concentration parameter of a Dirichlet process, which determines the tendency of data points to form distinct groups. The main difference between our model and theirs lies in the way the dependence is induced. While their models specify a dependence structure on the atoms within a hierarchical model on a product space, our model introduces dependence directly at the level of partitions, providing a more streamlined approach to modeling partition relationships.
Without loss of generality, we begin by examining the case $T=2$, where $T$ can now be interpreted as the number of different views, groups or experimental conditions, depending on the context of the application. Thus, we focus on the joint distribution of $\pi_{1}$ and $\pi_{2}$, as implied by the PSM, which is
@@
$\tilde{\pi}\sim p^{*}(\tilde{\pi}).$ (13)
@@
where $\tilde{\gamma}_{t}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{Bern}(\tilde{\eta})$, $t=1,2$, controls the whether $\pi_{t}$ coincides with the parent partition $\tilde{\pi}$. A higher value of $\tilde{\eta}$ increases the chances of $\pi_{t}$ being different from $\tilde{\pi}$, allowing for more flexibility and individual variations in the partition structure. Conversely, a lower value of $\tilde{\eta}$ makes it more likely that $\pi_{t}$ will follow the same partition structure as $\tilde{\pi}$, enforcing a stronger dependence between the two. Note that, in model (13)–(14), the distribution of $\pi_{2}$ is independent of that of $\pi_{1}$, given the partition $\tilde{\pi}$. The joint distribution of the random partitions $\pi_{1}$ and $\pi_{2}$, the common parent partition $\tilde{\pi}$, and the dependence variables $(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$, here denoted by $p(\bm{\pi}_{1:2},\tilde{\pi},\tilde{\gamma}_{1},\tilde{\gamma}_{2})$, can thus be written as
@@
The distribution of $\pi_{1}$ and $\pi_{2}$ is obtained by marginalizing the last expression with respect to $\tilde{\pi}$ and $(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$, which gives (12), provided that $\tilde{\eta}=1-\sqrt{1-\eta}$. See Section A.4 in the Supplementary Material. Hence, the local level partition model’s joint distribution on the partitions can be viewed as a special case of a hierarchical partition model, when considered marginally. This representation highlights the connection between the time-varying formulation of the PSM, as specified through (3)-(4), and an exchangeable model on the partitions, under the assumptions of identical marginal distributions of the partitions, $p^{*}(\pi_{t})$, and equal Bernoulli probabilities $\tilde{\eta}$ across groups. This hierarchical representation can be extended to accommodate $T\geq 2$ groups, with one parent random partition driving the distribution of $T$ identically distributed random partitions. Moreover, the assumption of a common Bernoulli probability $\tilde{\eta}$ across groups can be relaxed to accommodate for group-specific parameters $\tilde{\eta}_{t}$, with $t=1,\ldots,T$. Expressing the local level partition model as a hierarchical model provides an alternative viewpoint to appreciate its flexibility in capturing both similarities and dissimilarities among the partitions. The hierarchical structure allows for the sharing of information across groups while still accommodating individual variations in the partition structures.